### v2

# Enhancement of superconductivity on thin film of Sn under high pressure

Abstract
We investigated the pressure effects of a superconductivity on thin films of Sn. Elemental superconductor Sn with a body-centered tetragonal structure, -Sn, exhibits superconductivity below the superconducting transition temperature ( K) at ambient pressure. of Sn increases with lowering dimension such as in thin film and nanowire growth, or by high-pressure application. For thin films, exhibits a slight increase up to approximately 4 K compared to the bulk value, attributable to the crystalline size and lattice disorder. By applying pressure on a bulk Sn, initially decreases from 3.72 K as the pressure increases. Further increasing pressure up to 10 GPa, increases to 5.3 K with the structural transformation. However, the combination of these effects on thin films of Sn, namely, thin-film growth and pressure effects, remains underexplored. In this study, we combined film-growth and pressure-application techniques to further increase using a diamond anvil cell with boron-doped diamond electrodes. The drop of the electrical resistance suggesting the onset of on the thin film reached above 6 K in -Sn phase. Further, the upper critical magnetic field was drastically enhanced. Atomic force microscopy suggests that the refinement of the grain size of the thin film under the non-hydrostatic pressure conditions contributes to stabilizing the higher of -Sn.
I Introduction
Applying high pressures to directly compress a material is a useful approach to investigate intriguing physical properties and search for new materials Mao et al. (2018); Yamanaka (2010). For instance, oxygen—a gas at ambient conditions— exhibits metallic behaviors and superconductivity at high pressures Shimizu et al. (1998, 2005). Recently, high-temperature superconductors such as hydrogen-rich materials and nickelates under high pressure have attracted considerable attention Drozdov et al. (2015, 2019); Somayazulu et al. (2019); Matsumoto et al. (2020); Semenok et al. (2020); Chen et al. (2021); Kong et al. (2021); Troyan et al. (2021); Li et al. (2022); Ma et al. (2022); Song et al. (2023); Cross et al. (2024); Sun et al. (2023); Sakakibara et al. (2024). In addition, recent discoveries of the high-temperature superconducting states in elemental solid Ti and Sc at extreme pressures imply the potential of high-temperature superconductivity in high-pressure phases of other elements Liu et al. (2022); Zhang et al. (2022); Ying et al. (2023).
Some elemental superconductors with thin-film dimensions show an increase in superconducting transition temperature () compared to the bulk value, attributable to the crystalline size and lattice disorder Buckel and Hilsch (1954); Abeles et al. (1966); Garland et al. (1968); Strongin et al. (1968). In the case of Sn, for thin films and nano-wires varies slightly depending on the size and surface morphology Tian et al. (2005); Beutel et al. (2016); Bang et al. (2019); Lozano et al. (2019). The mechanism of enhancement is not thoroughly clear although it has been proposed to arise from changes in the phonon density of states, the electron density of states, and the electron-phonon coupling Knorr and Barth (1970); Houben et al. (2017, 2020).
Sn exhibits various crystal structures at high pressures Barnett et al. (2004); Salamat et al. (2013); Deffrennes et al. (2022). Half a century ago, Wittig investigated the electrical transport properties of bulk Sn at high pressures of up to 16 GPa and revealed the highest is 5.3 K at 11.3 GPa on the pressure induced phase, -Sn, where -Sn shows superconductivity below 3.72 K at ambient pressure Wittig (1966). However, research on combination of pressure application and thin-film growth on elemental superconductors is inadequate. Here, we hypothesize that changing the crystalline size of a thin film via pressure application could stabilize the higher .
In this study, we combined thin film growth and pressure-application techniques to increase using a diamond anvil cell (DAC) with boron-doped diamond (BDD) electrodes Matsumoto et al. (2016, 2017, 2018). We investigated the pressure effects on superconductivity of thin films of Sn compared to the bulk sample. We observed a higher for the thin film comapred to previous studies on high-pressure phase. Further, we observed a drastically enhanced critical magnetic field on the thin film under high pressure.
II Experimental procedure
Thin films of Sn were deposited on a diamond anvil by a resistance heating evaporation. The target metal was a high-purity Sn, purchased from Kojundo Chemical Lab. Co. Ltd. Comparing with the high pressure measurements, we also prepared Sn thin film on a diamond substrate for electrical transport measurements at ambient pressure. The film deposition on two diamonds was performed simultaneously. The optical images of the thin films on the diamond anvil and the diamond substrate are shown in Fig. 1(a) and the inset of Fig. 1(c), respectively. The film thickness and surface morphology of the films were evaluated via atomic force microscopy (AFM; Nanocute, SII NanoTechnology Inc.) at room temperature. For the magnetic measurements of the bulk Sn, a wire of high-purity Sn (Nilaco Corp.) was used.
High-pressure generation was performed using DAC. The pressure value at room temperature was evaluated using ruby fluorescence and Raman shift of diamond Mao et al. (1978); Akahama and Kawamura (2004). In the electrical transport measurements, diamond anvil with BDD electrodes with a culet size of 300 m was used Matsumoto et al. (2016, 2017, 2018). The BDD electrodes were deposited homoepitaxially on the diamond anvil by microwave plasma chemical vapor deposition Takano et al. (2004). This electrode exhibits high durability and can be reused until the diamond anvil itself fractures. Further, thin films could be deposited directly on the diamond anvil with BDD electrodes, eliminating the need for an electrode fabrication process after thin film deposition Matsumoto et al. (2016); Adachi et al. (2020); Matsumoto et al. (2021). The pressure-transmitting media in the solid, liquid, and gaseous states are compatible with this system. A gasket of a stainless steel was pre-indented and a 200 m diameter hole was drilled. The insulating layer was prepared using a MgO–epoxy mixture. We termed this setup as non-hydrostatic. We also performed the high-pressure generation with better hydrostatic pressure condition than non-hydrostatic measurement using a liquid pressure-transmitting medium, glycerol. A 150 m diameter hole was drilled in the insulating layer of MgO–epoxy mixture to prepare the sample space, which was filled with glycerol. We termed this setup as quasi-hydrostatic. In quasi-hydrostatic pressure measurements, the pressure value was evaluated using Raman shift of diamond Akahama and Kawamura (2004). The electrical transports under a magnetic field () perpendicular to the surface were measured by a four-terminal method by physical properties measurement system (Quantum Design).
For magnetic measurements, we used a miniature DAC in combination with a superconducting quantum interference device magnetometer (MPMS, Quantum Design) Mito et al. (2001, 2014, 2016, 2017); Abdel-Hafiez et al. (2018); Mito et al. (2019). A nano-polycrystalline diamond with culet size of 600 m and a pre-indented tungsten gasket with a hole size of 200 m were used Irifune et al. (2003). Bulk Sn pieces and ruby powders were loaded into the sample space without a pressure-transmitting medium. The in-phase component of the AC magnetic response () was measured. The frequency and amplitude of the AC field were 3 Hz and 0.2 mT, respectively.
III Experimental Results
III.1 Ambient pressure
Figure 1(b) shows the temperature () dependence of for the bulk Sn under , where no background signal from DAC is subtracted. Below 3.7 K, the diamagnetic signal suggesting the superconducting state was observed at . The onset was decreased by applying . Figure 1(c) shows the dependence of the electrical resistance () under perpendicular to the film surface. The residual resistance ratio (RRR) was estimated to be 11. At , an drop suggesting the onset of the superconducting transition was observed at 3.75 K. slightly increased compared to the bulk value and was similar to values from the previous studies on thin films Beutel et al. (2016); Bang et al. (2019). By applying , the onset of decreased with increasing . However, the critical magnetic field was three times higher than that of the bulk Sn (Fig. 1(d)). of 100 mT estimated using was similar to that of a previous study Bang et al. (2019). Considering to the , the thin film transformed into a type II superconductor, as previously reported Dolan and Silcox (1973); Bang et al. (2019); Ohkuma et al. (2023).
III.2 Non-hydrostatic pressure on bulk Sn
Figure 2(a) shows the dependence of on the bulk Sn under high pressures. On applying pressure, decreased, as previously reported. Figure 2(b) shows the dependence of at 2.7 GPa under varying . The decrease of was observed by applying mT. At mT, was below 2 K. was evaluated to be 13 mT, as shown in the inset of Fig. 2(b).
III.3 Non-hydrostatic pressure on thin film
Figure 3(a) shows the dependence of on the thin film of Sn under high pressures on pressurization. As shown in the inset of Fig. 4(a), RRR was estimated to be 1.4 at 2.5 GPa, which was much lower than that under ambient pressure. We speculate that the crystallographic defect was introduced because of the non-hydrostatic pressure condition. The drop suggesting the superconducting transition was observed around 4 K, whereas was 3.75 K under ambient pressure. decreased on further increasing the pressure, as previously reported. showed peak behavior just above at 5.5 GPa, possibly due to the granularity or disorder on the thin films. Above 9.5 GPa, slightly tended to decrease around 6 K, suggesting the superconducting transition, where -Sn phase could emerge. With further pressure application, the drop became significant, and slightly decreased. After applying 20 GPa, the pressure was decreased to 10.5 GPa. Figure 3(b) shows the dependence of for the thin film of Sn under high pressures with depressurization. The onset of increased to 6.3 K at 10.5 GPa, and -Sn remained at 8.5 GPa. -Sn vanished when the pressure was decreased to 3.5 GPa.
Figures 4(a) and (b) show the dependence of for the thin film of Sn under various at 2.5 and 10.5 GPa under non-hydrostatic pressure condition. was observed at 1 T, whereas the upper critical magnetic field () was approximately 0.1 T under ambient pressure. Drastic enhancement was also observed in the phase at 10.5 GPa. As shown in Fig. 4(c), estimated using Werthamer–Helfand–Hohenberg (WHH) model reached several teslas at high pressures Helfand and Werthamer (1966); Werthamer et al. (1966); Baumgartner et al. (2014). Notably, the enhancement of -Sn under the non-hydrostatic pressure condition was observed reproducibly. Figure 5(b) shows the dependence of for the thin film of Sn with the other setup at 5 GPa. The optical image of the thin film of Sn is shown in Fig. 5(a). The decrease suggesting the superconducting transition was observed even at T.
There are two possible scenarios for the enhancement. The first is the enhancement of the flux-pinning force. The crystallographic defects introduced by non-hydrostatic pressure could serve as pinning centers, potentially leading to an increase in . The other is the shortening of the mean free path of the electrons and the resulting shortening of the coherence length. Considering that the thickness of the thin film was approximately 100 nm, the electrical resistivity of the normal state near was estimated to be approximately for ambient pressure measurement. This value is consistent with those reported in previous studies Bang et al. (2019); Ohkuma et al. (2023). In contrast, the electrical resistivities of the thin films for 2.5 GPa of Run 1 and 5 GPa of Run 2 were approximately and , respectively. The increase in the electrical resistivity near suggests that the mean free path of the electrons is shortened by application of pressure.
III.4 Quasi-hydrostatic pressure on thin film
We performed high-pressure generation with better hydrostatic pressure condition using a liquid pressure transmitted medium, glycerol. Figure 6(a) shows the – of the thin film at 9 GPa under quasi-hydrostatic pressure conditions. The optical image of the thin film inside the DAC is shown in the inset of Fig. 6(a). The RRR was with in 2–3 in quasi-hydrostatic pressure measurements, which was slightly higher than those for the non-hydrostatic pressure condition. The drop suggesting superconducting transition was observed around 5 K. Figure 6(b) shows – under quasi-hydrostatic pressure condition between 2 and 8 K with warming process. The value was normalized using the value at 8 K. At 9 GPa, slightly decreased with decreasing around 5.3 K, suggesting the superconducting transition on -Sn. The drop was also observed around 3 K. The -Sn phase became significant on further pressure application. The onset of slightly increased at 12 GPa. We also decreased the pressure from 12 to 8 GPa. The decrease suggesting the superconducting transition of -Sn was slightly observed, whereas the -Sn was clearly observed at 8.5 GPa under the non-hydrostatic pressure condition. With further pressure decrease, -Sn vanished at 2 GPa.
IV Discussion
IV.1 Pressure dependence of
Figure 7 shows the pressure dependence of for the thin film and bulk samples compared with results from a previous study Wittig (1966). For the bulk Sn, the behavior of with respect to pressure was good agreement with the previous study. A similar tendency was observed for the thin film in -Sn phas;, however, its was higher compared to bulk Sn. One possible reason is the geometry of the thin film. As shown in Fig 1(a), the thin film area occupies approximately 70 of the culet of the diamond anvil, which produces the pressure distribution. In the -Sn phase, with quasi-hydrostatic pressure measurements showed a trend similar to the previous results Wittig (1966). Meanwhile, with non-hydrostatic pressure showed a higher value. We observed the highest of 6.3 K at 10.5 GPa, which was approximately 10 higher than that reported in a previous study Wittig (1966). The highest is not fully explained by the pressure gradient. Assuming that the of bulk -Sn continues to change at a rate of K/GPa, it is necessary to decrease the pressure to 4 GPa from the phase above 10 GPa. However, the phase cannot exist metastably under this pressure as was observed during the to phase transition on the decompression process (Fig. 3(b)). We emphasize that we observed the mid point of to be greater than 6.0 K (Fig. 4(b)).
IV.2 Atomic force microscopy
To investigate the morphology on the thin films, we performed AFM measurements before and after applying pressure under the non-hydrostatic pressure condition. Figures 8(a) and (b) show the optical and AFM images of the thin film before pressurization. The thin film of Sn was deposited on a diamond anvil without BDD electrodes. The average film thickness was evaluated to be 70 nm. A magnified AFM image is shown in Fig. 8(c). The average grain size was estimated to be approximately 300 nm. Figure 8(d) shows the optical image of the thin film after applying pressure up to 4.2 GPa. The pressure was evaluated by the diamond Raman shift. Most of the thin film peeled off and was transferred to the MgO–epoxy mixture. We measured the small remaining thin-film area. Figure 8(e) shows the AFM image of the thin film after pressurization. The average film thickness was evaluated to be 60 nm. A magnified view is shown in Fig. 8(f). The grain size was reduced to several tens of nanometers on pressure application.
Next, we performed AFM measurements under quasi-hydrostatic pressure condition for comparison with the non-hydrostatic pressure measurements. The thin film was the same as that used in the – measurement. Figures 9(a) and (b) show the AFM images of the thin film before applying pressure. The average thickness and average grain size were evaluated to be 82 and 600 nm, respectively. Figures 9(c) and (d) show the AFM images on the thin film after applying pressure up to 12 GPa. The pressure-transmitting medium was removed using ethanol and nitrogen gas. The average thickness was evaluated to be 91 nm. Unlike the non-hydrostatic pressure measurements, grain refinement was not observed.
The AFM results revealed that the grain refinement was observed only under the non-hydrostatic pressure condition. In the – results, the maximum value under the non-hydrostatic condition was 10 higher than that under the quasi-hydrostatic condition, suggesting that the grain refinement plays a pivotal role in stabilizing the higher . Smaller grain sizes tend to have higher , because of phonon softening under ambient pressure Houben et al. (2020). Houben performed nuclear resonant inelastic x-ray scattering on nano-structured films and bulk Sn to investigate the phonon density of states and observed a decrease in the high-energy phonon modes and a slight increase in the low-energy phonon modes in nano-structured films Houben et al. (2017, 2020). Using the obtained phonon spectra, calculations based on the Allen–Dynes–McMillan formalism yielded values in good agreement with the experimental data. In nano-structured films, the electron–phonon coupling increased by up to 10, suggesting that phonon softening and the associated change in electron–phonon coupling play a major role in the increase. We consider that the increase under the non-hydrostatic pressure condition is related to the grain size and that grain refinement could induce to changes in electron-phonon coupling.
IV.3 Possible higher on Sn
Recently, anomalies in magnetization, resistance, and heat capacity suggesting the superconducting transition were observed around 5.5 K in nano-wires of Sn Zhang et al. (2016). Further, based on scanning tunneling microscopy, the thin film of Sn deposited on substrate exhibited superconductivity around 8 K Shao et al. (2018). Indeed, in our thin film, the decrease of was observed at 11 K under 9.5 GPa, suggesting the signature of the superconducting transition, and the anomaly shifted to lower by applying the magnetic field (Fig 10). On the other hand, some granular or amorphous thin films exhibit a resistance decrease at higher than due to the effects of fluctuations Glover (1967); Aslamasov and Larkin (1968); Skocpol and Tinkham (1975); Sóñora et al. (2019). Further investigations such as magnetic measurements Hsieh et al. (2019); Lesik et al. (2019), heat capacity measurements Umeo (2016), and scanning tunneling microscopy Cao et al. (2023) under high pressure may offer insights for the possible stabilization of higher .
V Conclusion
In conclusion, we demonstrated the pressure effect on the superconductivity of thin films of Sn. We observed the superconductivity below 6.3 K in the -phase of Sn, which was approximately higher than previous bulk results. Further, the drastically increased under the non-hydrostatic high pressure condition. We also observed the signature of the superconducting transition at higher than . AFM results suggest that the grain refinement under non-hydrostatic pressure contributes to the stabilization of the higher of -Sn.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant No. 19H02177, 20H05644 and JST-Mirai Program Grant No. JPMJMI17A2. This work was also supported by the Joint Usage/Research Center PRIUS, Ehime University, JapanData Availability Statement
The data that support the findings of this article are openly available Sasaki et al. (2025)